# AI for the Open-World: the Learning Principles

### v1

@@ 4.2.4.1 Comparison of two approaches
To answer the question above, this section compares the performance (open-world AI ability) [^16] In i.i.d.regime, such as persistent-knowledge storing and retrieval, of course, more data + larger model = better performance. This argument in i.i.d. scenario has been well studied three decades ago [Vapnik 1991a]. of Memory Mosaics v2 and Transformer Large models trained on different amounts of data. Specifically, multiple Transformer Large models are trained on 200b, 1t, and 8t training tokens, while a Memory Mosaics v2 Large model is trained on 1t training tokens.
@@
Figures 4.10 and 4.11 show the comparison on in-context learning tasks. ×$\times 8$ times more training data helps Transformer Large (8t data) match the performance of Memory Mosaics v2 (1t data) on semantic label tasks (Figure 4.10). However, more training data cannot help Transformer Large on the more difficult anonymous label tasks. In contrast, more training data (8t) hurts Transformer Large on anonymous label tasks (Figure 4.11).
@@ 4.3 Discussion
Future works of inference-time learning.
@@
Currently in Memory Mosaics v2, the hyperparameter of the memory-based method (i.e. bandwidth of Gaussian kernel smoothing) is determined by the “input statistics” (e.g. number of examples) instead of cross-validation. It is well known that cross-validation is (almost) always the best way to choose hyperparameters. Automatically selecting bandwidth parameter at inference-time using “cross-validation” is another appealing future direction.
@@ A.2.1 Experiment settings
A.2.3 Experiments on a deeper architecture: resnet152
@@
A.2.4 Fine-tuning experiments
@@
A.2.5 Vision transformer Experiment settings
@@ Appendix C Inference-time learning
C.1 Memory Mosaics V2 with long-term memory removed after training
Table C.1 shows removing long-term memory from Memory Mosaics V2 Large during inference degrades the performance on ruler question-answer tasks by 20$\sim$30%. It indicates that the ruler question-answer tasks rely on long-term memory to perform well.

### v2

@@ 4.2.4.1 Comparison of two approaches
To answer the question above, this section compares the new tasks learning ability (open-world AI ability) [^19] In i.i.d.regime, such as persistent-knowledge storing and retrieval, of course, more data + larger model = better performance. This argument in i.i.d. scenario has been well studied three decades ago [Vapnik 1991a]. of Memory Mosaics v2 and Transformer Large models trained on different amounts of data. Specifically, multiple Transformer Large models are trained on 200b, 1t, and 8t training tokens, while a Memory Mosaics v2 Large model is trained on 1t training tokens.
New-knowledge storage and retrieval
@@
In-context learning
Figures 4.8 and 4.9 show the comparison on in-context learning tasks. ×$\times 8$ times more training data helps Transformer Large (8t data) match the performance of Memory Mosaics v2 (1t data) on semantic label tasks (Figure 4.8). However, more training data cannot help Transformer Large on the more difficult anonymous label tasks. In contrast, more training data (8t) hurts Transformer Large on anonymous label tasks (Figure 4.9).
@@ 4.3 Discussion
4.3.2 Future works
@@
It is well known that the reconstruction-based objective function is helpful for the compact language data. However, for other completed data, such as video, early work [Lerer et al. 2016] reveals the weakness of this reconstruction-based objective function. That is, the prediction tends to be “blur”. Incorporating an objective function working on latent space (rather than input space), such as Joint Embedding [LeCun 2022], into Memory Mosaics v2 is another interesting research direction.
Currently in Memory Mosaics v2, the hyperparameter of the memory-based method (i.e. bandwidth of Gaussian kernel smoothing) is determined by the “input statistics” (e.g. number of examples) instead of cross-validation. It is well known that cross-validation is (almost) always the best way to choose hyperparameters. Automatically selecting bandwidth parameter, as well as some other essential parameters, at inference-time using “cross-validation” is another appealing future direction.
@@ A.2.1 Experiment settings
A.2.2 Experiments on a deeper architecture: resnet152
@@
A.2.3 Fine-tuning experiments
@@
A.2.4 Vision transformer Experiment settings
@@ Appendix C Inference-time learning
C.1 Additional results for new-knowledge storage and retrieval
Table C.1 shows that removing long-term memory from memory mosaics v2 after training degrades the performance on the ruler question-answer tasks by 20$\sim$30%. This indicates that the ruler question-answer tasks rely on long-term memory to perform well.
@@ C.1 Additional results for new-knowledge storage and retrieval
Table C.2 compares memory mosaics v2 large and other public base models on ruler question-answer tasks. Memory mosaics v2 large outperforms these models across all task lengths.
Table C.2: Comparison of Memory Mosaics v2 large (base model) and other public base models (similar scale) on ruler question-answer tasks. Memory Mosaics v2 large outperforms these models across all task lengths. The numbers in “*” rows come from Hsieh et al. [2024].
Model claimed length task-length 4k 8k 16k 32k
Memory-Mosaics-v2-large (base) 32k 58.9 55.5 54.9 53.4
Llama2-7B (base)* 4k 48.6 - - -
Mixtral-base (8x7B)* 32k 50.8 47.7 45.3 41.3
Mistral-base (7B)* 32k 53.5 51.0 48.4 44.7
Together-base (7B)* 32k 47.5 44.6 33.6 0.0
LongLoRA-base (7B)* 100k 34.5 32.1 33.6 29.4
Yarn-base (7B)* 128k 29.7 23.5 28.6 29.7
LWM-base (7B)* 1M 42.7 40.2 38.7 37.1